# CRADLE Bench: A Clinician-Annotated Benchmark for Multi-Faceted Mental Health Crisis and Safety Risk Detection

### v1

@@ 2 Related Work
The CLPsych 2019 shared task (Zirikly et al., 2019) introduced a Reddit-based benchmark with expert-derived risk detection guidelines, later extended by the IEEE BigData 2024 Cup (Li et al., 2024) into a multi-class task. Nguyen and Pham (2024) used LLMs to generate pseudo-labels, showing that weak supervision can enhance suicide ideation detection. SHINES dataset and CESM-100 dataset (Ghosh et al., 2025) improved LLMs’ ability to distinguish casual mentions from serious self-harm intent. Beyond suicide risk, the SafeCity corpus (Karlekar and Bansal, 2018) contains real-world reports of sexual harassment, supporting multi-label classification of harassment types. The Webis Trigger Warning Corpus (Wiegmann et al., 2023) provides fanfiction documents with up to 36 trigger warnings, while the X-Sensitive dataset (Antypas et al., 2025) integrates six categories such as profanity, sexually explicit material, or drug related content from Twitter. PsyEvent dataset (Lv et al., 2025) captures life events in Reddit posts and shows its utility for suicide risk prediction. Deng et al. (2025) introduced PsyCrisisBench, a dataset of annotated Chinese psychological hotline transcripts labeled for mood, suicidal ideation, suicide planning, and risk assessment, while Gaur et al. (2019) leveraged the Columbia-Suicide Severity Rating Scale (C-SSRS; Posner et al. 2011) to propose a five-level suicide risk classification for more fine-grained prediction.
@@ 3.3.4 Temporal Dimensions of Crises
Table 3: Distribution of temporal labels in the development and test sets. Percentages are computed relative to the total number of labels in each split.
@@ 4.1 Models
Multiple LLMs are tested on CRADLE BENCH, including Llama (Grattafiori et al., 2024), Gemma (Team et al., 2025), Qwen (Qwen et al., 2025; Yang et al., 2025), gpt-oss (OpenAI et al., 2025), Gemini-2.5-Pro (Comanici et al., 2025), Claude-4-Sonnet (Anthropic, 2025), and GPT-5 (OpenAI, 2025).
@@ 4.2 Results
Table 4 presents the performance of different models on CRADLE Bench. In general, larger models tend to outperform their smaller counterparts across most metrics. However, there are exceptions: for example, gemma-3-12B performs better than gemma-3-27B on several metrics, likely due to increased false positives from more liberal label predictions. Within the open-source families, Llama-3.3-70B and Qwen2.5-72B achieve strong results, with Llama reaching a Jaccard score of 0.7806 and a Micro F1 of 0.7773. The best-performing open-source model is gpt-oss-120b, which attains the highest Exact Match score of 0.7650 and competitive performance across other metrics.
@@ 5.2 Training Details
We fine-tune three open-source LLMs as base models: Qwen3-14B, chosen for its competitive performance relative to model size. We also select Llama-3.3-70B-Instruct and Qwen2.5-72B-Instruct, which demonstrate strong overall results on CRADLE BENCH. All models are fine-tuned on the training sets described above. Detailed training configurations and hyperparameters are provided in Appendix F.

### v2

@@ 2 Related Work
The CLPsych 2019 shared task (Zirikly et al., 2019) introduced a Reddit-based benchmark with expert-derived risk detection guidelines, later extended by the IEEE BigData 2024 Cup (Li et al., 2024) into a multi-class task. Nguyen and Pham (2024) used LLMs to generate pseudo-labels, showing that weak supervision can enhance suicide ideation detection. SHINES dataset and CESM-100 dataset (Ghosh et al., 2025) improved LLMs’ ability to distinguish casual mentions from serious self-harm intent. Beyond suicide risk, the SafeCity corpus (Karlekar and Bansal, 2018) contains real-world reports of sexual harassment, supporting multi-label classification of harassment types. The Webis Trigger Warning Corpus (Wiegmann et al., 2023) provides fanfiction documents with up to 36 trigger warnings, while the X-Sensitive dataset (Antypas et al., 2025) integrates six categories such as profanity, sexually explicit material, or drug related content from Twitter. PsyGuard (Qiu et al., 2024) is an automated framework for identifying suicidal ideation and assessing risk in psychological counseling settings. PsyEvent dataset (Lv et al., 2025) captures life events in Reddit posts and shows its utility for suicide risk prediction. Deng et al. (2025) introduced PsyCrisisBench, a dataset of annotated Chinese psychological hotline transcripts labeled for mood, suicidal ideation, suicide planning, and risk assessment, while Gaur et al. (2019) leveraged the Columbia-Suicide Severity Rating Scale (C-SSRS; Posner et al. 2011) to propose a five-level suicide risk classification for more fine-grained prediction.
@@ 3.3.4 Temporal Dimensions of Crises
Table 3: Distribution of temporal labels in the development and test sets. Each crisis label is assigned exactly one temporal marker (ongoing, past, or no_crisis). Percentages are computed as (count of labels with that temporal marker) / (total number of labels in split), thus summing to 100%.
@@ 4.1 Models
Multiple state-of-the-art LLMs are tested on CRADLE BENCH, including Llama (Grattafiori et al., 2024), Gemma (Team et al., 2025), Qwen (Qwen et al., 2025; Yang et al., 2025), gpt-oss (OpenAI et al., 2025), Gemini-2.5-Pro (Comanici et al., 2025), Claude-4-Sonnet (Anthropic, 2025), and GPT-5 (OpenAI, 2025).
@@ 4.2 Results
Table 4 presents the performance of different models on CRADLE Bench. In general, larger models tend to outperform their smaller counterparts across most metrics. However, there are exceptions: for example, gemma-3-12B performs better than gemma-3-27B on several metrics, likely due to increased false positives from more liberal label predictions. Within the open-source families, Llama-3.3-70B and Qwen2.5-72B achieve strong results, with Llama reaching a Jaccard score of 0.7806 and a Micro F1 of 0.7773. The best-performing open-source model is gpt-oss-120b, which attains the highest Exact Match score of 0.7650, Jaccard of 0.7996, and competitive performance across other metrics.
@@ 5.2 Training Details
We fine-tune three open-source LLMs as base models: Qwen3-14B, chosen for its competitive performance relative to model size. We also select Llama-3.3-70B-Instruct and Qwen2.5-72B-Instruct, which demonstrate strong overall results on CRADLE BENCH. All models are fine-tuned on the training sets described above. Detailed training configurations and hyperparameters are provided in Appendix F. Fine-tuned models are publicly released (links in Appendix F).
@@ 7 Conclusion
Acknowledgments
We gratefully acknowledge the support of the Doo- Good Foundation. Any opinions, findings, and con- clusions or recommendations expressed in this ma- terial are those of the authors and do not necessarily reflect the views of the DooGood Foundation.
We would also like to thank Ashley Stinson, PhD for her assistance with annotation for this study.
@@ Appendix F Training Setup
Fine-tuned Models
Models can be found here: Model Links Qwen3-14B consensus, unanimous Llama-3.3-70B-Instruct consensus, unanimous Qwen2.5-72B-Instruct consensus, unanimous